Ahmer walks 795m north and 55m west 80m south find the displacement and distance?
1
Expert's answer
2020-03-06T10:26:53-0500

The distance


"d=795+55+80=930\:\rm m."

The displacement


"s=\sqrt{(795-80)^2+55^2}=717\:\rm m."